OpenStudy (anonymous):

In ∆ABC, AB = 4.12, BC = 4, AC = 5. What is m C?

OpenStudy (anonymous):

you got a picture? looks like this should be a law of cosines problem does that seem right?

OpenStudy (anonymous):

Law of cosines is correct

OpenStudy (anonymous):

no theres no pictures and yes thats what it is

OpenStudy (anonymous):

lets draw one and label it

OpenStudy (anonymous):

ok idk how to tho

OpenStudy (anonymous):



OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

my picture stinks, because 4 is larger than 5, but mo matter at least we have something to refer to

OpenStudy (anonymous):

\[c^2=a^2+b^2-2ab\cos(C)\] we are looking for \(C\)

OpenStudy (anonymous):

no its fine

OpenStudy (anonymous):

solving for \(\cos(C)\) you get \[\cos(C)=\frac{a^2+b^2-c^2}{2ab}\]

OpenStudy (anonymous):

now we put in the numbers you are given

OpenStudy (anonymous):

\[\cos(C)=\frac{4^2+5^2-(4.12)^2}{2\times 4\times 5}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and so \(C\) is the arccosine of that number we need a calculator to find it

OpenStudy (anonymous):

i just plugged it in directly to wolfram and it gave me this http://www.wolframalpha.com/input/?i=arccos%28%285^2%2B4^2-4.12^2%29%2F40%29

OpenStudy (anonymous):

\[\frac{ 16+25-(4.12)^2 }{40 ? }\]

OpenStudy (anonymous):

yes that is the cosine of the angle you want the angle itself so you want \[\cos^{-1}\left(\frac{ 16+25-(4.12)^2 }{40 }\right)\]

OpenStudy (anonymous):

typed it in to wolfram and got about \(53.08\)